Find the principal needed now to get the given amount, that is, find the present value.
To get $3000 after 33/4 years at 5% compounded daily
The present value of $3000 is $?

Answer by mathmate(429)

Question:
Find the principal needed now to get the given amount, that is, find the present value.
To get $3000 after 33/4 years at 5% compounded daily

Solution:
F=3000
APR=5%
i=0.05/365 (daily interest rate)
n=33/4*365 (days)
Use compound interest formula
F=P(1+i)^n
3000=P(1+0.05/365)^(33/4*365)
=>
P=3000/(1+0.05/365)^(33/4*365)
=3000/1.51055
=$1986.04 [present value required to accumulate 3000 at APR=5% for 33/8 years]